One number is 8 more than another. Their sum is 94. Find the numbers.
Let the numbers = x
and their sum = 94
the numbers are?

Found 2 solutions by ewatrrr, lawdej:
Answer by ewatrrr(24785)

 
Hi,
One number is 8 more than another
Let x and (x+8) represent the numbers
Question states*** Their sum is 94.
x + (x+8) = 94
2x+8 = 94
2x = 86
x = 43
the numbers are 43 and 53 (43+8)
CHECKING our Answer***
43 + 51 = 94

Answer by lawdej(58)

Let the numbers x and y
x+y=94 and x=8+y
8+y+y=94
2y=94-8
2y=86
y=86/2
y=43
x=8+y=8+43=51
so the numbers are 43 and 51
(lawdej@yahoo.com)